OpenStudy (anonymous):

Find the equation of the line through the point (–2, –2) that is perpendicular to the line y = 2x – 1.

OpenStudy (wach):

Ok, so first, what constitutes a perpendicular line? The slope has the to be the inverse of the line, given, right?

OpenStudy (wach):

So what would the inverse of 2x as a slope be ?

OpenStudy (wach):

It would be -1/2 right? We basically take 2/1 and flip it over, and then change the sign. So this one is a complete inverse. For the point it runs through, we can use point slope formula: (y+2)=-1/2(x+2)

OpenStudy (wach):

From there, you can leave the line in that form, or simplify for y=mx+b

OpenStudy (wach):

I hope that helps!